Alice needs to make a certain number of boxes. If she makes the boxes at a rate of 16 boxes per hour, she will take 2 hours longer than if she works at a rate of 18 boxes per hour. How many boxes does Alice need to make?
Answer by stanbon(75887)

Alice needs to make a certain number of boxes. If she makes the boxes at a rate of 16 boxes per hour, she will take 2 hours longer than if she works at a rate of 18 boxes per hour. How many boxes does Alice need to make?
------------------
Original DATA:
# of boxes needed:"x";rate=16boxes/hr;time=x boxes/16 boxes/hr = x/16 hrs
------
New DATA:
# of boxes needed:"x";rate=18boxes/hr;time=x boxes/18 boxes/rs = x/18 hrs
---
Equation:
old time - new time = 2 hrs
x/16 - x/18 = 2
---
18x - 16x = 2*16*18
2x = 2*16*18
x = 16*18 = 288 boxes (# of boxes Alice needs to make)
=================
Cheers,
Stan H.
===============